# A spin on Hagedorn temperatures and string stars

### v1

@@ 2.3 A Euclidean derivation of $\beta_{H}(\nu)$
Following the work of Sathiapalan:1986db ; Atick:1988si , there is a second, intuitive way to explain the Hagedorn instability of string theory in the canonical ensemble. The idea is to follow Polchinski:1985zf , and understand the canonical partition function $Z(\beta)$ as the Euclidean string theory partition function on $\mathbb{R}^{d}\times S^{1}_{\beta}$ (with anti-periodic boundary condition for the fermions). In terms of the Euclidean theory, the Hagedorn instability is a divergence of the 1-loop path integral in this background. Unlike field theory, string theory in this background includes sectors of strings winding around the thermal circle. Due to their extension along the thermal circle, the mass of the winding strings is, up to an overall constant, proportional to $m^{2}(\beta)\sim\beta^{2}$. A direct computation shows that the lightest winding mode turns massless at the Hagedorn temperature $\beta_{H}$, and tachyonic for higher temperatures. As it becomes massless, the 1-loop partition function acquires a new divergence at $\beta=\beta_{H}$, and turns ill-defined for $\beta<\beta_{H}$. This is the Euclidean reasoning behind the linear Hagedorn growth of the entropy.
@@
Following Polchinski:1985zf , the trace (16) is simply equal to the string theory partition function on the background (22) under the identification (23).
Treating $X^{0}$ as a spatial circle, these spaces are known as Melvin spaces. While originally proposed as solutions of Einstein-Maxwell theory Melvin:1963qx , they were later embedded in KK theory Gibbons:1987ps ; Dowker:1993bt . The KK reduction of (22) includes a non-trivial (but asymptotically flat) curved metric, dilaton, and KK magnetic flux profile localized around the center of the rotation plane $\Phi=0$ (also known as a “magnetic fluxbrane”). As an exact Lorentzian string theory model, it was discussed, for example, in  Tseytlin:1994ei ; Dowker:1995sg ; Russo:1995tj ; Gutperle:2001mb . [^9] We thank Roberto Emparan for introducing us to this subject. To the best of our knowledge, the following thermal interpretation we suggest is novel. However, it is interesting to note that the possible relation between this background and a rotating black hole was already discussed in Dowker:1995gb .
@@
We use light-cone quantization, choosing two of the other d2d-2 directions for gauge-fixing Polchinski:1998rq . To find the spectrum of the winding sector, we need to determine the worldsheet ground-state energy in that sector. The contribution of $X^{0}$ and all the other d2d-2 coordinates to the ground-state energy is standard. The complex scalar $\Phi$, however, has twisted boundary conditions which shift the standard harmonics from $\omega_{n}=|n|$ to $\omega_{n}=|n+\nu|$. As a result, the $\Phi$ contribution to the ground-state energy is [^10] Using the identity Polchinski:1998rq $\sum_{n=1}^{\infty}(n+\nu)=-\frac{1}{12}-\frac{1}{2}(\nu+\nu^{2}).$ (25)
@@
To find the mass of the lightest winding mode, we need to ask what the lowest allowed level $N$ is, as a function of $\nu$. At $\nu=0$Atick:1988si showed that we should take $N=0$. At $\nu=1$, however, the target-space fermions are periodic on the thermal circle. Using T-duality, the small $R$ limit is dual to the decompactified $\mathbb{R}^{10}$ limit. In this limit, we know that the tachyon $N=0$ is projected out by GSO, and the lightest state has $N=1/2$ Polchinski:1998rr . What is the correct value for $N$? Another way to phrase the question is the following. Although the full theory is expected to go back to itself only under $\nu\sim\nu+2$, the mass-shell condition (31) appears to have a shorter periodicity of $\nu\sim\nu+1$, as both (26) and (30) have that periodicity. How is that possible?
@@
As we show in appendix A, one can obtain the same Hagedorn growth from direct state counting. At $\nu=0$, we get back the standard answer $R_{H}=\sqrt{2}l_{s}$ (6Atick:1988si . As in the bosonic case, we find that increasing $\nu$ initially increases the Hagedorn temperature to leading order. This time, however, we find that at $\nu=1$ the Hagedorn temperature is infinity ($R_{H}=0$)! We can understand this result as follows. At $\nu=1$, the target space has periodic boundary conditions for the fermions. In the $R\ll l_{s}$ limit, this mode is T-dual to a graviton momentum mode of $g_{00}$. We expect such a mode to turn massless $m^{2}=0$ in the non-compact $R=0$ limit, which gives $R_{H}=0$ in this case. In fact, due to the periodic boundary conditions, at $\nu=1$ we have $e^{2\pi i\nu J}=(-1)^{F}$ (where $F$ is the target space fermion number), and the partition function actually computes the supersymmetric index of the theory Urbach:2025eyu . Due to cancellations inside the multiplets, we don’t expect the computation to depend on the temperature, and specifically, there is no Hagedorn instability. As a result, (34) should strictly hold only for $\nu\neq 1\text{ mod }2$.
Finally, we turn to heterotic strings. At $\nu=0$, the GSO projection dictates that the momentum along $X_{0}$ is an odd half-integer (instead of an integer) Atick:1988si . To generalize the analysis to $\nu\neq 0$, we keep the momentum $n$ unfixed and write down the right and left mass shell conditions. For the right-movers, we have the twisted boson (27) together with $32$ untwisted fermions
@@
$n=\frac{1+\nu^{2}}{2},\quad\alpha^{\prime}M^{2}=\frac{R^{2}}{\alpha^{\prime}}+\frac{\alpha^{\prime}}{R^{2}}\frac{(1+\nu^{2})^{2}}{4}+\nu^{2}+2|\nu|-3.$ (38)
@@
For a direct state-counting derivation of this Hagedorn temperature, see appendix A. As a sanity check, note that this result coincides with the standard $R_{H}/l_{s}=1+1/\sqrt{2}$ at $\nu=0$ Gross:1985fr . As in the type II case, there is no Hagedorn divergence or instability at $\nu=1$ due to supersymmetric cancellations, and (39) should be strictly understood only for $\nu\neq 1\text{ mod }2$. [^11] Note that here, unlike type II, we have a finite “Hagedorn temperature” $R_{H}=l_{s}$ at $\nu=1$. The behavior of the winding mode and the string star solution close to $\nu=1$ will be discussed elsewhere Urbach:2025eyu .
@@ 2.4 A target space analysis
We would like to explain the results of the previous section also from a target space perspective. At $\nu=0$, the target space perspective was put forward by Atick and Witten Atick:1988si . Reducing on the thermal $S^{1}_{R}$ for $R-R_{H}\ll l_{s}$, the Euclidean string theory includes a light particle associated with the (lightest) winding mode, with the $d$-dimensional quadratic action
@@ 3 The (slowly) rotating string star
In the previous section, we studied properties of free strings. In this section, we turn on the string coupling $g_{s}$, or the gravitational constant $G_{N}$. Instead of the free string partition function (16), we will consider the string theory partition function with only asymptotically $\mathbb{R}^{d}\times S^{1}_{R}$ geometry (times some compact manifold to saturate the appropriate critical dimension, which we will suppress), twisted by some $\nu\neq 0$ (23). Perturbatively in $G_{N}$, we can consider different saddles of the partition function with these boundary conditions. One such saddle is the twisted $\mathbb{R}^{d}\times S^{1}_{R}$ itself, representing weakly-interacting thermal strings in flat space. Another is the (rotating) Euclidean black hole saddle, in which the thermal circle shrinks with an $S^{d-1}\times\mathbb{R}^{2}$ topology. This saddle remains under control (weakly-curved) only for $\beta\gg l_{s}$. At $\nu=0$ and for temperatures close to the Hagdorn temperature, Horowitz:1997jc found another type of saddle, a ‘string star’. Using the EFT found in Atick:1988si coupled to gravity, the authors found a normalizable condensate of the winding mode $\chi$ for $3\leq d\leq 5$ - a string star. In this section, we will find a similar saddle for the case of $\nu\neq 0$, a rotating string star.
Analogously to the $\nu=0$ case, we work at temperatures close to the Hagedorn temperature $R_{H}(\nu)$. As the winding mode becomes parametrically light compared to the other stringy modes, we can look for saddles of the EFT derived in the previous section, together with gravity. Analogously to the nonrotating case Chen:2021dsw ; Horowitz:1997jc , we will keep only quadratic fluctuations of the metric and a cubic interaction of the metric components with the winding mode, and verify a posteriori that the solution is self-consistent. Due to the more complicated structure of the metric, more metric components couple at this order to the winding mode compared to the $\nu=0$ case, including also the dilaton.
@@ 3.1 The effective action
The low-energy effective action is composed of two parts $I=I_{\text{grav}}+I_{\chi}$. $I_{\text{grav}}$ is the (super)gravity low-energy action of the string theory in question, around $\mathbb{R}^{d}\times S^{1}$. As we are going to excite only the winding mode $\chi$, it is enough to consider the backreaction of the metric and the dilaton Horowitz:1997jc , with the action [^14] At higher orders in $R-R_{H}\ll l_{s}$, the winding mode could also excite a non-zero NS-NS flux.
@@
To leading order in $R-R_{H}\ll l_{s}$, it is enough to expand the gravitational action $I_{\text{grav}}$ to quadratic order, and study the cubic interaction of $\chi$ through $I_{\chi}$ Chen:2021dsw ; Urbach:2022xzw . It is useful to introduce a $d$-dimensional dilaton field
@@
where, following Chen:2021dsw , we have introduced the notation
@@ 3.2 Perturbative solutions
First, arbitrarily close to the modified Hagedorn temperature $(R-R_{H}(\nu))/l_{s}\ll\nu\ll 1$, we have $m_{0}^{2}\alpha^{\prime}\sim-|\nu|$. In this regime, we expect the solution profile to be similar to the winding mode solution. Therefore, we expect (see (51)) $L_{||}\sim l_{s}/\sqrt{\nu}$ while $L_{\perp}\sim l_{s}^{\frac{3}{2}}/\sqrt{R-R_{H}(\nu)}$. In other words, the solution extends mostly in the perpendicular direction and remains localized in the rotation plane. [^15] For $d=2$, the equations for $\varphi$ and $\chi$ are similar to the $AdS_{3}$ string star equations Urbach:2023npi at leading order in $l_{s}/l_{ads}$, with $|\nu|=R_{H}/l_{ads}$. The analogy to the $AdS_{3}$ string star equations means that there is a solution to these equations in $d=2$. However, there is no ‘nonrotating’ limit for such a solution, just as there’s no flat space limit for the $AdS_{3}$ string star. See section 5 for more information about the situation in AdS.
@@
Lastly, we can consider $\nu$ sufficiently small so that we can treat the problem perturbatively in $\nu^{2}$, and solve the equations order by order. At $\nu=0$, the spherically-symmetric solution has a typical length scale of $L\sim 1/m_{0}$ Chen:2021dsw ; Horowitz:1997jc . Since the leading $\nu$ interaction term is controlled by $\nu^{2}\rho^{2}/\alpha^{\prime}$, demanding it to be perturbative requires that $\nu^{2}L^{2}\ll R^{2}-R_{H}^{2}$. In terms of $m_{0}^{2}$, that means [^16] Since the EFT is valid only for $\alpha^{\prime}m_{0}^{2}\ll 1$, this also entails $\nu\ll 1$ (62).
@@
To write this equation, we moved to the $r,\theta$ variables introduced in (52), and took $\chi$ to be real. As these are the $\nu=0$ equations, the first two equations are the standard (non-rotating) Horowitz-Polchinski equations Horowitz:1997jc ; Chen:2021dsw . By absorbing $m_{0}$ into a redefinition of $r$, the equations are scale-invariant. As a result, the solution’s length scale and amplitude are $\sim 1/m_{0}$ and $\alpha^{\prime}m_{0}^{2}$ respectively. Together, this justifies the gradient expansion and quadratic approximation we made a posteriori, in the near-Hagedorn temperature limit $m_{0}^{2}\alpha^{\prime}\ll 1$. The same holds for the novel $\Psi_{0}$ equation. Normalizable solutions to the equations (66) exist (numerically) for $3\leq d\leq 5$, see figure 2. For $d\geq 6$, there are no self-consistent solutions for the EFT, as the solution will always have an instability towards shrinking down Chen:2021dsw ; Horowitz:1997jc (see Balthazar:2022szl ; Balthazar:2022hno ; Bedroya:2024igb ; Chu:2025boe ; Chu:2025kzl for recent progress about stringy solutions at $d>5$).

### v2

@@ 2.3 A Euclidean derivation of $\beta_{H}(\nu)$
Following the work of Sathiapalan (1987); Kogan (1987); Atick and Witten (1988), there is a second, intuitive way to explain the Hagedorn instability of string theory in the canonical ensemble. The idea is to follow Polchinski (1986), and understand the canonical partition function $Z(\beta)$ as the Euclidean string theory partition function on $\mathbb{R}^{d}\times S^{1}_{\beta}$ (with anti-periodic boundary condition for the fermions). In terms of the Euclidean theory, the Hagedorn instability is a divergence of the 1-loop path integral in this background. Unlike field theory, string theory in this background includes sectors of strings winding around the thermal circle. Due to their extension along the thermal circle, the mass of the winding strings is, up to an overall constant, proportional to $m^{2}(\beta)\sim\beta^{2}$. A direct computation shows that the lightest winding mode turns massless at the Hagedorn temperature $\beta_{H}$, and tachyonic for higher temperatures. As it becomes massless, the 1-loop partition function acquires a new divergence at $\beta=\beta_{H}$, and turns ill-defined for $\beta<\beta_{H}$. This is the Euclidean reasoning behind the linear Hagedorn growth of the entropy.
@@
Following Polchinski (1986), the trace (16) is simply equal to the string theory partition function on the background (22) under the identification (23).
Treating $X^{0}$ as a spatial circle, these spaces are known as Melvin spaces. While originally proposed as solutions of Einstein-Maxwell theory Melvin (1964), they were later embedded in KK theory Gibbons and Maeda (1988); Dowker et al. (1994). The KK reduction of (22) includes a non-trivial (but asymptotically flat) curved metric, dilaton, and KK magnetic flux profile localized around the center of the rotation plane $\Phi=0$ (also known as a “magnetic fluxbrane”). As an exact Lorentzian string theory model, it was discussed, for example, in  Tseytlin (1995); Dowker et al. (1996); Russo and Tseytlin (1995); Gutperle and Strominger (2001). [^9] We thank Roberto Emparan for introducing us to this subject. To the best of our knowledge, the following thermal interpretation we suggest is novel. However, it is interesting to note that the possible relation between this background and a rotating black hole was already discussed in Dowker et al. (1995).
@@
We use light-cone quantization, choosing two of the other d2d-2 directions for gauge-fixing Polchinski (2007a). To find the spectrum of the winding sector, we need to determine the worldsheet ground-state energy in that sector. The contribution of $X^{0}$ and all the other d2d-2 coordinates to the ground-state energy is standard. The complex scalar $\Phi$, however, has twisted boundary conditions which shift the standard harmonics from $\omega_{n}=|n|$ to $\omega_{n}=|n+\nu|$. As a result, the $\Phi$ contribution to the ground-state energy is [^10] Using the identity Polchinski (2007a) $\sum_{n=1}^{\infty}(n+\nu)=-\frac{1}{12}-\frac{1}{2}(\nu+\nu^{2}).$ (25)
@@
To find the mass of the lightest winding mode, we need to ask what the lowest allowed level $N$ is, as a function of $\nu$. At $\nu=0$Atick and Witten (1988) showed that we should take $N=0$. At $\nu=1$, however, the target-space fermions are periodic on the thermal circle. Using T-duality, the small $R$ limit is dual to the decompactified $\mathbb{R}^{10}$ limit. In this limit, we know that the tachyon $N=0$ is projected out by GSO, and the lightest state has $N=1/2$ Polchinski (2007b). What is the correct value for $N$? Another way to phrase the question is the following. Although the full theory is expected to go back to itself only under $\nu\sim\nu+2$, the mass-shell condition (31) appears to have a shorter periodicity of $\nu\sim\nu+1$, as both (26) and (30) have that periodicity. How is that possible?
@@
As we show in appendix A, one can obtain the same Hagedorn growth from direct state counting. At $\nu=0$, we get back the standard answer $R_{H}=\sqrt{2}l_{s}$ (6Atick and Witten (1988). As in the bosonic case, we find that increasing $\nu$ initially increases the Hagedorn temperature to leading order. This time, however, we find that at $\nu=1$ the Hagedorn temperature is infinity ($R_{H}=0$)! We can understand this result as follows. At $\nu=1$, the target space has periodic boundary conditions for the fermions. In the $R\ll l_{s}$ limit, this mode is T-dual to a graviton momentum mode of $g_{00}$. We expect such a mode to turn massless $m^{2}=0$ in the non-compact $R=0$ limit, which gives $R_{H}=0$ in this case. In fact, due to the periodic boundary conditions, at $\nu=1$ we have $e^{2\pi i\nu J}=(-1)^{F}$ (where $F$ is the target space fermion number), and the partition function actually computes the supersymmetric index of the theory Urbach (2025). Due to cancellations inside the multiplets, we don’t expect the computation to depend on the temperature, and specifically, there is no Hagedorn instability. As a result, (34) should strictly hold only for $\nu\neq 1\text{ mod }2$.
Finally, we turn to heterotic strings. At $\nu=0$, the GSO projection dictates that the momentum along $X_{0}$ is an odd half-integer (instead of an integer) Atick and Witten (1988). To generalize the analysis to $\nu\neq 0$, we keep the momentum $n$ unfixed and write down the right and left mass shell conditions. For the right-movers, we have the twisted boson (27) together with $32$ untwisted fermions
@@
$n=\frac{1+\nu^{2}}{2},\quad\alpha^{\prime}M^{2}=\frac{R^{2}}{\alpha^{\prime}}+\frac{\alpha^{\prime}}{R^{2}}\frac{(1+\nu^{2})^{2}}{4}-\nu^{2}+2|\nu|-3.$ (38)
@@
For a direct state-counting derivation of this Hagedorn temperature, see appendix A. As a sanity check, note that this result coincides with the standard $R_{H}/l_{s}=1+1/\sqrt{2}$ at $\nu=0$ Gross et al. (1985). As in the type II case, there is no Hagedorn divergence or instability at $\nu=1$ due to supersymmetric cancellations, and (39) should be strictly understood only for $\nu\neq 1\text{ mod }2$. [^11] Note that here, unlike type II, we have a finite “Hagedorn temperature” $R_{H}=l_{s}$ at $\nu=1$. The behavior of the winding mode and the string star solution close to $\nu=1$ will be discussed elsewhere Urbach (2025).
@@ 2.4 A target space analysis
We would like to explain the results of the previous section also from a target space perspective. At $\nu=0$, the target space perspective was put forward by Atick and Witten Atick and Witten (1988). Reducing on the thermal $S^{1}_{R}$ for $R-R_{H}\ll l_{s}$, the Euclidean string theory includes a light particle associated with the (lightest) winding mode, with the $d$-dimensional quadratic action
@@ 3 The (slowly) rotating string star
In the previous section, we studied properties of free strings. In this section, we turn on the string coupling $g_{s}$, or the gravitational constant $G_{N}$. Instead of the free string partition function (16), we will consider the string theory partition function with only asymptotically $\mathbb{R}^{d}\times S^{1}_{R}$ geometry (times some compact manifold to saturate the appropriate critical dimension, which we will suppress), twisted by some $\nu\neq 0$ (23). Perturbatively in $G_{N}$, we can consider different saddles of the partition function with these boundary conditions. One such saddle is the twisted $\mathbb{R}^{d}\times S^{1}_{R}$ itself, representing weakly-interacting thermal strings in flat space. Another is the (rotating) Euclidean black hole saddle, in which the thermal circle shrinks with an $S^{d-1}\times\mathbb{R}^{2}$ topology. This saddle remains under control (weakly-curved) only for $\beta\gg l_{s}$. At $\nu=0$ and for temperatures close to the Hagdorn temperature, Horowitz and Polchinski (1998) found another type of saddle, a ‘string star’. Using the EFT found in Atick and Witten (1988) coupled to gravity, the authors found a normalizable condensate of the winding mode $\chi$ for $3\leq d\leq 5$ - a string star. In this section, we will find a similar saddle for the case of $\nu\neq 0$, a rotating string star.
Analogously to the $\nu=0$ case, we work at temperatures close to the Hagedorn temperature $R_{H}(\nu)$. As the winding mode becomes parametrically light compared to the other stringy modes, we can look for saddles of the EFT derived in the previous section, together with gravity. Analogously to the nonrotating case Chen et al. (2023); Horowitz and Polchinski (1998), we will keep only quadratic fluctuations of the metric and a cubic interaction of the metric components with the winding mode, and verify a posteriori that the solution is self-consistent. Due to the more complicated structure of the metric, more metric components couple at this order to the winding mode compared to the $\nu=0$ case, including also the dilaton.
@@ 3.1 The effective action
The low-energy effective action is composed of two parts $I=I_{\text{grav}}+I_{\chi}$. $I_{\text{grav}}$ is the (super)gravity low-energy action of the string theory in question, around $\mathbb{R}^{d}\times S^{1}$. As we are going to excite only the winding mode $\chi$, it is enough to consider the backreaction of the metric and the dilaton Horowitz and Polchinski (1998), with the action [^14] At higher orders in $R-R_{H}\ll l_{s}$, the winding mode could also excite a non-zero NS-NS flux.
@@
To leading order in $R-R_{H}\ll l_{s}$, it is enough to expand the gravitational action $I_{\text{grav}}$ to quadratic order, and study the cubic interaction of $\chi$ through $I_{\chi}$ Chen et al. (2023); Urbach (2022). It is useful to introduce a $d$-dimensional dilaton field
@@
where, following Chen et al. (2023), we have introduced the notation
@@ 3.2 Perturbative solutions
First, arbitrarily close to the modified Hagedorn temperature $(R-R_{H}(\nu))/l_{s}\ll\nu\ll 1$, we have $m_{0}^{2}\alpha^{\prime}\sim-|\nu|$. In this regime, we expect the solution profile to be similar to the winding mode solution. Therefore, we expect (see (51)) $L_{||}\sim l_{s}/\sqrt{\nu}$ while $L_{\perp}\sim l_{s}^{\frac{3}{2}}/\sqrt{R-R_{H}(\nu)}$. In other words, the solution extends mostly in the perpendicular direction and remains localized in the rotation plane Urbach (2025).  [^15] For $d=2$, the equations for $\varphi$ and $\chi$ are similar to the $AdS_{3}$ string star equations Urbach (2023) at leading order in $l_{s}/l_{ads}$, with $|\nu|=R_{H}/l_{ads}$. The analogy to the $AdS_{3}$ string star equations means that there is a solution to these equations in $d=2$. However, there is no ‘nonrotating’ limit for such a solution, just as there’s no flat space limit for the $AdS_{3}$ string star. See section 5 for more information about the situation in AdS.
@@
Lastly, we can consider $\nu$ sufficiently small so that we can treat the problem perturbatively in $\nu^{2}$, and solve the equations order by order. At $\nu=0$, the spherically-symmetric solution has a typical length scale of $L\sim 1/m_{0}$ Chen et al. (2023); Horowitz and Polchinski (1998). Since the leading $\nu$ interaction term is controlled by $\nu^{2}\rho^{2}/\alpha^{\prime}$, demanding it to be perturbative requires that $\nu^{2}L^{2}\ll R^{2}-R_{H}^{2}$. In terms of $m_{0}^{2}$, that means [^16] Since the EFT is valid only for $\alpha^{\prime}m_{0}^{2}\ll 1$, this also entails $\nu\ll 1$ (62).
@@
To write this equation, we moved to the $r,\theta$ variables introduced in (52), and took $\chi$ to be real. As these are the $\nu=0$ equations, the first two equations are the standard (non-rotating) Horowitz-Polchinski equations Horowitz and Polchinski (1998); Chen et al. (2023). By absorbing $m_{0}$ into a redefinition of $r$, the equations are scale-invariant. As a result, the solution’s length scale and amplitude are $\sim 1/m_{0}$ and $\alpha^{\prime}m_{0}^{2}$ respectively. Together, this justifies the gradient expansion and quadratic approximation we made a posteriori, in the near-Hagedorn temperature limit $m_{0}^{2}\alpha^{\prime}\ll 1$. The same holds for the novel $\Psi_{0}$ equation. Normalizable solutions to the equations (66) exist (numerically) for $3\leq d\leq 5$, see figure 2. For $d\geq 6$, there are no self-consistent solutions for the EFT, as the solution will always have an instability towards shrinking down Chen et al. (2023); Horowitz and Polchinski (1998) (see Balthazar et al. (2022, 2024); Bedroya and Wu (2025); Chu and Kutasov (2025a, b) for recent progress about stringy solutions at $d>5$).